# Rua Dr. Bento Teobaldo Ferraz 271, 01140-070, São Paulo - SP, Brasil

### v1

@@ Abstract
Not much is known about superstring scattering amplitudes in curved backgrounds. Using the hybrid formalism in $\rm AdS_{3}\times S^{3}$ with pure NS-NS three-form flux, we compute a $\rm PSU(1,1|2)$-covariant three-point amplitude for half-BPS vertex operators inserted in the $\rm AdS_{3}$ boundary and show that it agrees with the RNS computation. The zero-mode prescription for the fermions in $\rm AdS$ is defined in terms of the “standard” spacetime SUSY generator. It is found that integrating out the fermionic worldsheet fields in the path integral gives rise to the target-space vielbein, which explicitly encodes that the conformal group on the boundary is identified with the symmetry group of the $\rm AdS$ bulk.
@@ 1 Introduction
Establishing a thorough understanding of covariant superstring formulations in Anti de-Sitter (AdS) backgrounds is a much needed step for a deeper comprehension of the dualities connecting the superstring with conformal field theories (CFT’s) living on the boundary of the AdS spacetime. Even though a handful of covariant sigma-model actions are explicitly known Berkovits:1999xv Berkovits:1999zq Berkovits:2004xu , vertex operators and superstring amplitude computations in $\rm AdS$ are widely unexplored Berkovits:2022ivl .
@@
In this work, we will make use of the pure NS-NS hybrid formalism in $\rm AdS_{3}\times S^{3}$ and compute a manifestly $\rm PSU(1,1|2)$-covariant three-point amplitude for half-BPS vertex operators inserted at a position $\mathbf{x}$ in the $\rm AdS_{3}$ boundary. This will be done after defining curved worldsheet fields by making use of the vielbein field
@@
Since spacetime supersymmetric superstring scattering amplitudes in curved backgrounds have been hardly ever investigated, this construction can have some important applications. In the first place, it provides intuition for what happens after one integrate out the worldsheet fermions in a general $\rm AdS$ background amplitude computation. Secondly, it gives insights about what the correct amplitude prescription in the more interesting case of $\rm AdS_{5}\times S^{5}$ target-space might be. There have been significant works over the last years on trying to understand superstring vertex operators Berkovits:2019rwq Fleury:2021ieo , and the correct amplitude measure for the fermionic fields $\theta^{\alpha}$ in the $\rm AdS_{5}\times S^{5}$ pure spinor formalism Berkovits:2008ga Berkovits:2019ulm .
@@
Vertex operators for the pure R-R and pure NS-NS $\rm AdS_{3}$ backgrounds were studied in refs. Dolan:1999dc and Langham:2000ac in the hybrid formalism. In addition, for the pure R-R flux case, a three-graviton amplitude in the $\rm AdS_{3}$ bulk was calculated in ref. Bobkov:2002bx . For $k=1$ units of NS-NS flux, or tensionless point, superstring correlation functions for vertex operators inserted in the $\rm AdS_{3}$ boundary were computed in Dei:2020zui Dei:2023ivl and shown to match with the dual two-dimensional CFT Eberhardt:2018ouy . Moreover, using a near boundary limit — which is exact at $k=1$ — these correlators were also analyzed in refs. Knighton:2023mhq Knighton:2024qxd for bosonic strings and in ref. Sriprachyakul:2024gyl for the Ramond-Neveu-Schwarz (RNS) superstring.
@@
The content of this paper is the following. In Section 2, we review the hybrid formalism in $\rm AdS_{3}\times S^{3}$ background with pure NS-NS three-form flux together with spelling out our conventions for the worldsheet theory. In Section 3, we solve the physical state conditions and define half-BPS vertex operators in terms of a fermionic zero-mode coordinate $\theta^{\alpha}$. In Section 4, we perform a similarity transformation along the boundary direction and define the worldsheet fields and vertex operators depending on $\mathbf{x}\in{\rm\partial AdS_{3}}$ while introducing the vielbein field ${E_{A}}^{B}(\mathbf{x})$. In Section 5, we compute a $\rm PSU(1,1|2)$-covariant three-point amplitude for vertex operators inserted in the $\rm AdS_{3}$ boundary and show that integrating out the fermionic fields $\theta^{\alpha}$ imply the appearance of ${E_{A}}^{B}(\mathbf{x})$ in the kinematic factor. In Section 6, we further confirm our results by comparing with the RNS formalism and draw our conclusion in Section 7. At last, there exist a couple of appendices responsible for giving additional details.
@@ 2.1 Worldsheet action
Note that in the equation above we anti-symmetrize without dividing by the number of terms. In our notation, $\underline{a}=\{0$ to 5}5\}5 } is an $\rm SO(1,5)$ vector index, $\alpha=\{1$ to 4}4\}4 } is a fundamental $\rm SU(4)$ index and $j=\{1,2\}$ is an $\rm SU(2)$ index.
@@ 2.2 Superconformal generators
The $c=6$ $\mathcal{N}=2$ SCA (18) can be readily verified from the OPE’s (2) and (17). The more complicated properties to check come from the supercurrent $G^{+}$. As shown in ref. Berkovits:1999im , one way to fix the form of the superconformal generator $G^{+}$ is by demanding the naive generalization from flat to curved space to be invariant under the “non-standard” spacetime supersymmetries generated by $Q_{\alpha 2}$. Another involved consistency condition to prove of the algebra generators (18) is checking that the OPE of $G^{+}$ with itself is regular.
@@
which makes manifest its invariance under the non-standard SUSY’s generated by the charge
$Q_{\alpha 2}$ $=\oint\big{(}S_{\alpha 1}e^{-\rho-i\sigma}+S_{\alpha 2}\big{)}\,.$ (23)
@@
This notation is convenient, since half the spacetime supersymmetries in the hybrid superstring $Q_{\alpha j}$ act different than the zero modes of the ${\rm PSU(1,1|2)}_{k}$ SUSY currents $S_{\alpha j}$. The latter is a consequence of the presence of the $\{\rho,\sigma\}$-ghosts in the four SUSY’s $Q_{\alpha 2}$ of eq. (23).
@@ 2.3 Physical state conditions
The additional constraints $G^{-}_{0}\mathcal{V}=\widetilde{G}^{-}_{0}\mathcal{V}=0$ in eqs. (28) are convenient to further eliminate auxiliary degrees of freedom and imply a covariant gauge choice, e.g., they are equivalent to the Lorenz gauge condition for the open string sector Berkovits:1999im or to the Wess-Zumino gauge for compactifications of the superstring down to four dimensions Wess:1992cp Berkovits:1997zd Benakli:2021jxs . As we will see in Section 3.2, it is also possible to define a Wess-Zumino-type gauge in the $\rm AdS_{3}\times S^{3}$ hybrid formalism which turns out to be suited for performing amplitude computations.
@@
which corresponds to a vertex operator in the same picture as $\mathcal{V}$ when translating to the RNS language Berkovits:1999im . However, the important difference, when compared to the RNS formalism, is that $\mathcal{V}$ carries states both from the Ramond and Neveu-Schwarz sectors. In the RNS description, the Ramond states carry half-integer picture and the NS states carry integer picture. In fact, this is a crucial feature of the hybrid formalism. It treats Ramond and Neveu-Schwarz sectors in the same footing since it only uses worldsheet variables of integer conformal weight.
@@ 3.1 Superfields in $\rm AdS_{3}\times S^{3}$
Since the standard SUSY’s have the simple form (24), similar as in flat space, it is convenient to define $\mathcal{V}$ to be a superfield expanded in terms of a fermionic coordinate $\theta^{\alpha}$ which transforms as a Weyl spinor and is conjugate to $Q_{\alpha 1}$. More precisely, we define the superspace fermionic variable $\theta^{\alpha}$ by the property [^2] Recall that in our notation $Q_{\alpha 1}=\nabla_{\alpha 1}$, see eqs. (24) and (26).
@@ 3.2 Vertex operators for the massless states
For an object $X_{A}$ transforming in the representation $A$ of $\rm PSU(1,1|2)$, we define the covariant derivative as
@@ 4 Curved worldsheet fields
When writing a field $\mathcal{O}$ without any labels, it means that it only depends on the worldsheet coordinates $z$. Therefore, it is inserted in the position $\mathbf{x}=0$ in the boundary. [^3] I would like to thank Lucas Martins and Dennis Zavaleta for discussions regarding this point. For an operator function of any $\mathbf{x}\in\partial{\rm AdS}_{3}$, we will write $\mathcal{O}(\mathbf{x},z)$ — or simply $\mathcal{O}(\mathbf{x})$ — which is equivalent to $e^{\mathbf{x}\nabla_{+}}\mathcal{O}e^{-\mathbf{x}\nabla_{+}}$. As we will presently see, there are only a finite number of terms that contribute in this similarity transformation for our fundamental worldsheet variables.
@@ 4.2 Curved fermionic coordinates
As we discussed in the beginning of Section 3, the vertex $\mathcal{V}$ in eq. (47) is a superfield in our superstring description, i.e., it is a function of the fermionic zero-mode variable $\theta^{\alpha}$. In order to compute amplitudes involving the vertex $\mathcal{V}$ inserted in the $\rm AdS_{3}$ boundary, we need to specify what are the analogues of the superspace coordinates $\theta^{\alpha}$ when we introduce dependence on the spacetime coordinate $\mathbf{x}$.
@@
Therefore, from (35), one finds that the action of the standard SUSY’s on $\theta^{\alpha}(\mathbf{x})$ is
@@ 4.3 Some properties of the vielbein
$f_{\underline{a}\underline{b}\underline{c}}(\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3})$ $={E_{\underline{a}}}^{\underline{d}}(\mathbf{x}_{1}){E_{\underline{b}}}^{\underline{e}}(\mathbf{x}_{2}){E_{\underline{c}}}^{\underline{f}}(\mathbf{x}_{3})f_{def}\,,$ (74)
@@ 5 Amplitude computation
In Sections 3 and 4, after introducing the zero-mode fermionic coordinate $\theta^{\alpha}$, we determined the gauge-fixed vertex operators for the half-BPS states. Additionally, we showed how vertex operators inserted at $\mathbf{x}=0$ can be translated to an arbitrary position $\mathbf{x}$ in the $\rm AdS_{3}$ boundary by the means of a similarity transformation and using the vielbein ${E_{A}}^{B}(\mathbf{x})$.
@@ 5.2 Integrating out the fermions
where $D_{\underline{a}}(\mathbf{x}_{12})={E_{\underline{a}}}^{\underline{b}}(\mathbf{x}_{12})D_{\underline{b}}$ and $f_{\underline{a}\underline{b}\underline{c}}(\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3})$ are the covariant derivative (51) and the structure constants with curved indices (4.3). We are also using that
@@
The design of the amplitude (5.2) begs for an interpretation. As we have alluded to below eq. (73), the spacetime vielbein field ${E_{\underline{a}}}^{\underline{b}}(\mathbf{x})$ encodes that a conformal transformation in the $\rm AdS_{3}$ boundary corresponds to a rotation in the $\rm AdS_{3}$ bulk. In particular, this can be seen by the observation that the object ${E_{a}}^{+}\partial_{+}$ generates infinitesimal Möbius transformations along $\rm\partial AdS_{3}$.
@@ 6.1 From hybrid to RNS variables
Note that the RNS variables obey the usual OPE’s
$H_{C}^{\rm RNS}(y)H_{C}^{\rm RNS}(z)$ $\sim-2\log(y-z)\,,$ $\sigma(y)\sigma(z)$ $\sim-\log(y-z)\,.$ (87a)
@@ 6.3 Three-point amplitude in RNS variables
We can now use the tools developed in this section to compute the three-point amplitude (5.2) for the NS-sector states inserted in the $\rm AdS_{3}$ boundary directly in terms of the RNS formalism prescription.
@@
where ${E_{\underline{a}}}^{\underline{b}}(\mathbf{x})$ is given by (56). Hence, the fundamental OPE’s read
@@ 7 Conclusion
In this work, we have studied the superstring in the $\rm AdS_{3}\times S^{3}\times\mathcal{M}_{4}$ background (where $\mathcal{M}_{4}$ is a Calabi-Yau two-fold) with $k$ units of NS-NS self-dual three-form flux from the $\rm PSU(1,1|2)$ supergroup perspective. We used the Berkovits-Vafa-Witten hybrid formalism in $\rm AdS_{3}$ Berkovits:1999im and wrote a $\rm PSU(1,1|2)$-covariant three-point amplitude for half-BPS states inserted in the $\rm AdS_{3}$ boundary. As a corollary, we found that the kinematic factor gets dressed with the vielbein ${E_{A}}^{B}(\mathbf{x})$ after the worldsheet fermions are integrated out in the path-integral. In addition, we saw the compelling fact of the conformal group on the boundary being identified with the symmetry group of the $\rm AdS_{3}$ bulk by explicitly analyzing the form of ${E_{A}}^{B}(\mathbf{x})$, which naturally appears in our covariant superstring description. We ended by showing that our results agree with the RNS formalism answer.
Let us now comment on some related subjects that deserve future investigation. Although string theory scattering amplitudes in flat space have been extensively explored both at tree- and quantum-level Berkovits:2022ivl Mafra:2022wml , superstring amplitudes in curved backgrounds are still a mysterious subject. In view of the $\rm AdS/CFT$ correspondence, it is highly desirable to have superstring perturbation theory under good control in an $\rm AdS$ background. [^9] See ref. Alday:2023mvu for recent progress in the $\rm AdS_{5}\times S^{5}$ Virasoro-Shapiro amplitude. That being said, the construction presented in this paper provides all the necessary elements for the exploration of higher-point $\rm PSU(1,1|2)$-covariant tree-level scattering amplitudes in $\rm AdS_{3}\times S^{3}$.
@@
What happens when switching on R-R flux? Although the holomorphic/anti-holomorphic factorization is lost Benichou:2010rk Eberhardt:2018exh , the tree-level amplitude prescription to integrate over the fermions in Section 2.4 should not change when turning on a non-zero amount of R-R three-form flux Bobkov:2002bx . Given that the worldsheet action in a R-R $\rm AdS_{3}\times S^{3}$ background is known Berkovits:1999im , one could look for reproducing the results from refs. Cho:2018nfn Gaberdiel:2023lco using the spacetime supersymmetric hybrid formalism.
@@
So it seems reasonable to suspect that there is a prescription in the $\rm AdS_{5}\times S^{5}$ pure spinor formalism which picks out a factor of $(\theta)^{5}(\widehat{\theta})^{5}$. In the $\rm U(5)$ notation, by using the standard spacetime SUSY generator $\nabla_{a}$, the prescription should take the form
@@
where $a,\widehat{a}=1$ to $5$ are fundamental and anti-fundamental $\rm U(5)$ indices in (106). It would be interesting to first study these claims in the analogue of the $\rm AdS_{5}\times S^{5}$ pure spinor formalism in $\rm AdS_{3}\times S^{3}$ Berkovits:1999du CD25_2 by building up on the results of this paper. Note also that progress towards $\rm AdS_{5}\times S^{5}$ vertex operators was achieved in refs. Berkovits:2019rwq Fleury:2021ieo .

### v4

@@ Abstract
Not much is known about superstring scattering amplitudes in curved backgrounds. Using the hybrid formalism in $\rm AdS_{3}\times S^{3}$ with pure NS-NS three-form flux, we compute a $\rm PSU(1,1|2)$-covariant three-point amplitude for half-BPS vertex operators inserted on the $\rm AdS_{3}$ boundary and show that it agrees with the RNS computation. The zero-mode prescription for the fermions in $\rm AdS$ is defined in terms of the “standard” spacetime SUSY generator. It is found that integrating out the fermionic worldsheet fields in the path integral gives rise to the target-space vielbein, which explicitly encodes that the conformal group on the boundary is identified with the symmetry group of the $\rm AdS$ bulk.
@@ 1 Introduction
Establishing a thorough understanding of covariant superstring formulations in Anti de-Sitter (AdS) backgrounds is a much needed step for a deeper comprehension of the dualities connecting the superstring with conformal field theories (CFTs) living on the boundary of the AdS spacetime. Even though a handful of covariant sigma-model actions are explicitly known Berkovits:1999xv Berkovits:1999zq Berkovits:2004xu Metsaev:1998it Rahmfeld:1998zn Park:1998un , vertex operators and superstring amplitude computations in $\rm AdS$ are widely unexplored Berkovits:2022ivl .
@@
In this work, we will make use of the pure NS-NS hybrid formalism in $\rm AdS_{3}\times S^{3}$ with $k$ units of three-form flux and compute a three-point amplitude for half-BPS vertex operators inserted at a position $\mathbf{x}$ on the $\rm AdS_{3}$ boundary in a manifestly $\rm PSU(1,1|2)$-covariant fashion, i.e., using the spacetime supersymmetric worldsheet variables of the hybrid description. This will be done after defining curved worldsheet fields by making use of the vielbein field
@@
Since spacetime supersymmetric superstring scattering amplitudes in curved backgrounds have been hardly ever investigated, this construction can have some important applications. In the first place, it provides intuition for what happens after the worldsheet fermions are integrated out in a general $\rm AdS$ background amplitude computation. Secondly, it gives insights about what the correct amplitude prescription in the more interesting case of $\rm AdS_{5}\times S^{5}$ target-space might be. There have been significant works over the last years on trying to understand superstring vertex operators Berkovits:2019rwq Fleury:2021ieo , and the correct amplitude measure for the fermionic fields $\theta^{\alpha}$ in the $\rm AdS_{5}\times S^{5}$ pure spinor formalism Berkovits:2008ga Berkovits:2019ulm .
@@
Vertex operators for the pure R-R and pure NS-NS $\rm AdS_{3}$ backgrounds were studied in refs. Dolan:1999dc and Langham:2000ac in the hybrid formalism. In addition, for the pure R-R flux case, a three-graviton amplitude in the $\rm AdS_{3}$ bulk was calculated in ref. Bobkov:2002bx . For $k=1$ units of NS-NS flux, or tensionless point, superstring correlation functions for vertex operators inserted on the $\rm AdS_{3}$ boundary were computed in Dei:2020zui Dei:2023ivl and shown to match with the dual two-dimensional CFT Eberhardt:2018ouy . Moreover, using a near boundary limit — which is exact at $k=1$ — these correlators were also analyzed in refs. Knighton:2023mhq Knighton:2024qxd for the bosonic string and in ref. Sriprachyakul:2024gyl for the Ramond-Neveu-Schwarz (RNS) superstring.
There have been major advancements regarding string theory three-point functions in $\rm AdS_{3}$ with $k$ units of pure NS-NS flux and with vertex operators inserted in $\rm\partial AdS_{3}$. These advancements stem from both the RNS superstring formalism as well as from the bosonic $\rm SL(2,\mathbb{R})$ WZW model Maldacena:2001km Gaberdiel:2007vu Cardona:2009hk Teschner:1999ug Sriprachyakul:2024gyl . More recently, a closed formula, in terms of an integral expression, has been derived for three-point functions in the $\rm SL(2,\mathbb{R})$ WZW model involving arbitrary spectrally flowed states Dei:2021xgh Bufalini:2022toj .
The hybrid formalism for the superstring offers a significant advantage over the RNS formalism by enabling quantization while manifestly preserving some of the spacetime supersymmetry (SUSY) Berkovits:1996bf , and it eliminates the need for picture-changing or spin fields. Unlike the RNS formalism, the scattering amplitude prescription in the hybrid description is topological Berkovits:1994vy , so that there is no need for summing over spin structures as well. Moreover, as opposed to the Green-Schwarz action, the hybrid action is quadratic in a flat background.
@@
The content of this paper is the following. In Section 2, we review the hybrid formalism in $\rm AdS_{3}\times S^{3}$ background with pure NS-NS three-form flux together with spelling out our conventions for the worldsheet theory. In Section 3, we solve the physical state conditions and define half-BPS vertex operators in terms of a fermionic zero-mode coordinate $\theta^{\alpha}$. In Section 4, we perform a similarity transformation along the boundary direction and define the worldsheet fields and vertex operators depending on $\mathbf{x}\in{\rm\partial AdS_{3}}$ while introducing the vielbein field ${E_{A}}^{B}(\mathbf{x})$. In Section 5, we compute a $\rm PSU(1,1|2)$-covariant three-point amplitude for vertex operators inserted on the $\rm AdS_{3}$ boundary and show that integrating out the fermionic fields $\theta^{\alpha}$ imply the appearance of ${E_{A}}^{B}(\mathbf{x})$ in the kinematic factor. In Section 6, we further confirm our results by comparing with the RNS formalism, and draw our conclusion in Section 7. At last, there exist a couple of appendices responsible for giving additional details.
@@ 2.1 Worldsheet action
Note that in the equation above we anti-symmetrize with square brackets and without dividing by the number of terms. In our notation, $\underline{a}=\{0$ to 5}5\}5 } is an $\rm SO(1,5)$ vector index, $\alpha=\{1$ to 4}4\}4 } is a fundamental $\rm SU(4)$ index and $j=\{1,2\}$ is an $\rm SU(2)$ index.
@@ 2.2 Superconformal generators
The $c=6$ $\mathcal{N}=2$ SCA (18) can be readily verified from the OPEs (2) and (17). The more complicated properties to check come from the supercurrent $G^{+}$. As shown in ref. Berkovits:1999im , one way to fix the form of the superconformal generator $G^{+}$ is by demanding the naive generalization from flat to curved space to be invariant under the “non-standard” spacetime supersymmetries generated by $Q_{\alpha 2}$. Another involved consistency condition to prove of the algebra generators (18) is checking that the OPE of $G^{+}$ with itself is regular.
@@
which makes manifest its invariance under the non-standard SUSYs generated by the charge
$Q_{\alpha 2}$ $=\oint\big{(}S_{\alpha 1}e^{-\rho-i\sigma}+S_{\alpha 2}\big{)}\,,$ (23)
@@
This notation is convenient, since half the spacetime supersymmetries in the hybrid superstring $Q_{\alpha j}$ act different than the zero modes of the ${\rm PSU(1,1|2)}_{k}$ SUSY currents $S_{\alpha j}$. The latter is a consequence of the presence of the $\{\rho,\sigma\}$-ghosts in the four SUSYs $Q_{\alpha 2}$ of eq. (23).
@@ 2.3 Physical state conditions
The additional constraints $G^{-}_{0}\mathcal{V}=\widetilde{G}^{-}_{0}\mathcal{V}=0$ in eqs. (28) are convenient to further eliminate auxiliary degrees of freedom and imply a covariant gauge choice, e.g., they are equivalent to the Lorenz gauge condition for the open string sector Berkovits:1999im Wess:1992cp Berkovits:1997zd Benakli:2021jxs . As we will see in Section 3.2, it is also possible to define a Lorenz-type gauge in the $\rm AdS_{3}\times S^{3}$ hybrid formalism which turns out to be suited for performing amplitude computations.
@@
which corresponds to a vertex operator in the same picture as $\mathcal{V}$ when translating to the RNS language Berkovits:1999im . However, the important difference, when compared to the RNS formalism, is that $\mathcal{V}$ carries states both from the Ramond and Neveu-Schwarz sectors. In the RNS description, the Ramond states carry half-integer picture and the NS states carry integer picture. In fact, this is a crucial feature of the hybrid formalism. It treats Ramond and Neveu-Schwarz sectors in the same footing, since it only uses worldsheet variables of integer conformal weight.
@@ 3.1 Superfields in $\rm AdS_{3}\times S^{3}$
Since the standard SUSYs have the simple form (24), similar as in flat space, it is convenient to define $\mathcal{V}$ to be a superfield expanded in terms of a fermionic coordinate $\theta^{\alpha}$ which transforms as a Weyl spinor and is conjugate to $Q_{\alpha 1}$. More precisely, we define the superspace fermionic variable $\theta^{\alpha}$ by the property [^2] Recall that in our notation $Q_{\alpha 1}=\nabla_{\alpha 1}$, see eqs. (24) and (26).
@@ 3.2 Vertex operators for the massless states
For an object $X_{A}$ transforming in the representation $A$ of $\rm PSU(1,1|2)$, we define
@@ 4 Curved worldsheet fields
When writing a field $\mathcal{O}$ without any labels, it means that it only depends on the worldsheet coordinates $z$. Therefore, it is inserted in the position $\mathbf{x}=0$ on the boundary. [^3] I would like to thank Lucas Martins and Dennis Zavaleta for discussions regarding this point. For an operator function of any $\mathbf{x}\in\partial{\rm AdS}_{3}$, we will write $\mathcal{O}(\mathbf{x},z)$ — or simply $\mathcal{O}(\mathbf{x})$ — which is equivalent to $e^{\mathbf{x}\nabla_{+}}\mathcal{O}e^{-\mathbf{x}\nabla_{+}}$. As we will presently see, there are only a finite number of terms that contribute in this similarity transformation for our fundamental worldsheet variables.
@@ 4.2 Curved fermionic coordinates
As we discussed in the beginning of Section 3, the vertex $\mathcal{V}$ in eq. (47) is a superfield in our superstring description, i.e., it is a function of the fermionic zero-mode variable $\theta^{\alpha}$. In order to compute amplitudes involving the vertex $\mathcal{V}$ inserted on the $\rm AdS_{3}$ boundary, we need to specify what are the analogues of the superspace coordinates $\theta^{\alpha}$ when we introduce dependence on the spacetime coordinate $\mathbf{x}$.
@@
Therefore, from (35), one finds that the action of the standard SUSYs on $\theta^{\alpha}(\mathbf{x})$ is
@@ 4.3 Some properties of the vielbein
$f_{\underline{a}\underline{b}\underline{c}}(\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3})$ $={E_{\underline{a}}}^{\underline{d}}(\mathbf{x}_{1}){E_{\underline{b}}}^{\underline{e}}(\mathbf{x}_{2}){E_{\underline{c}}}^{\underline{f}}(\mathbf{x}_{3})f_{\underline{d}\underline{e}\underline{f}}\,,$ (74)
@@ 5 Amplitude computation
In Sections 3 and 4, after introducing the zero-mode fermionic coordinate $\theta^{\alpha}$, we determined the gauge-fixed vertex operators for the half-BPS states. Additionally, we showed how vertex operators inserted at $\mathbf{x}=0$ can be translated to an arbitrary position $\mathbf{x}$ on the $\rm AdS_{3}$ boundary by the means of a similarity transformation and using the vielbein ${E_{A}}^{B}(\mathbf{x})$.
@@ 5.2 Integrating out the fermions
where $D_{\underline{a}}(\mathbf{x}_{12})={E_{\underline{a}}}^{\underline{b}}(\mathbf{x}_{12})D_{\underline{b}}$ and $f_{\underline{a}\underline{b}\underline{c}}(\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3})$ are defined in eqs. (51) and (4.3), respectively. We are also using that
@@
The design of the amplitude (5.2) begs for an interpretation. As we have alluded to below eq. (73), the spacetime vielbein field ${E_{\underline{a}}}^{\underline{b}}(\mathbf{x})$ encodes that a conformal transformation on the $\rm AdS_{3}$ boundary corresponds to a rotation in the $\rm AdS_{3}$ bulk. In particular, this can be seen by the observation that the object ${E_{a}}^{+}\partial_{+}$ generates infinitesimal Möbius transformations along $\rm\partial AdS_{3}$.
@@ 6.1 From hybrid to RNS variables
Note that the RNS variables obey the usual OPEs
$H_{C}^{\rm RNS}(y)H_{C}^{\rm RNS}(z)$ $\sim-2\log(y-z)\,,$ $\sigma(y)\sigma(z)$ $\sim-\log(y-z)\,,$ (87a)
@@ 6.3 Three-point amplitude in RNS variables
We can now use the tools developed in this section to compute the three-point amplitude (5.2) for the NS-sector states inserted on the $\rm AdS_{3}$ boundary directly in terms of the RNS formalism prescription.
@@
where ${E_{\underline{a}}}^{\underline{b}}(\mathbf{x})$ is given by (56). Hence, the fundamental OPEs read
@@ 7 Conclusion
In this work, we have studied the superstring in the $\rm AdS_{3}\times S^{3}\times\mathcal{M}_{4}$ background (where $\mathcal{M}_{4}$ is a Calabi-Yau two-fold) with $k$ units of NS-NS self-dual three-form flux from the $\rm PSU(1,1|2)$ supergroup perspective. We used the Berkovits-Vafa-Witten hybrid formalism in $\rm AdS_{3}$ Berkovits:1999im and wrote a $\rm PSU(1,1|2)$-covariant three-point amplitude for half-BPS states inserted on the $\rm AdS_{3}$ boundary, whose coordinates are labelled by $\mathbf{x}$. As a corollary, we found that the kinematic factor gets dressed with the vielbein ${E_{A}}^{B}(\mathbf{x})$ after the worldsheet fermions are integrated out in the path-integral. In addition, we saw the compelling fact of the conformal group on the boundary being identified with the symmetry group of the $\rm AdS_{3}$ bulk by explicitly analyzing the form of ${E_{A}}^{B}(\mathbf{x})$, which naturally appears in our covariant superstring description. We ended by showing that our results agree with the RNS formalism answer.
Let us now comment on some related subjects that deserve future investigation. An immediate application of this work is the computation of the amplitude expression in components involving two R-sector states and one NS-sector state starting from the superspace form (5.1).
Although string theory scattering amplitudes in flat space have been extensively explored both at tree- and quantum-level Berkovits:2022ivl Mafra:2022wml , superstring amplitudes in curved backgrounds are still a mysterious subject. In view of the $\rm AdS/CFT$ correspondence, it is highly desirable to have superstring perturbation theory under good control in an $\rm AdS$ background. [^9] See ref. Alday:2023mvu for recent progress in the $\rm AdS_{5}\times S^{5}$ Virasoro-Shapiro amplitude. That being said, the construction presented in this paper provides all the necessary elements for the exploration of higher-point $\rm PSU(1,1|2)$-covariant tree-level scattering amplitudes in $\rm AdS_{3}\times S^{3}$.
@@
What happens when switching on R-R flux? Although the holomorphic/anti-holomorphic factorization is lost Benichou:2010rk Eberhardt:2018exh , the tree-level amplitude prescription to integrate over the fermions in Section 2.4 should not change when turning on a non-zero amount of R-R three-form flux Bobkov:2002bx . Given that the worldsheet action in a R-R $\rm AdS_{3}\times S^{3}$ background is known Berkovits:1999im , one could look for reproducing the results from refs. Cho:2018nfn Gaberdiel:2023lco Frolov:2023pjw using the spacetime supersymmetric hybrid formalism.
@@
So it seems reasonable to suspect that there is a prescription in the $\rm AdS_{5}\times S^{5}$ pure spinor formalism which picks out a factor of $(\theta)^{5}(\widehat{\theta})^{5}$. In the $\rm U(5)$ notation, by using the standard spacetime SUSY generator $\nabla_{a}$, the prescription may take the form
@@
where $a,\widehat{a}=1$ to $5$ are fundamental and anti-fundamental $\rm U(5)$ indices in (106). It would be interesting to first study these claims in the analogue of the $\rm AdS_{5}\times S^{5}$ pure spinor formalism in $\rm AdS_{3}\times S^{3}$ Berkovits:1999du Daniel:2024yfr by building up on the results of this paper. Note also that progress towards $\rm AdS_{5}\times S^{5}$ vertex operators was achieved in refs. Berkovits:2019rwq Fleury:2021ieo .